Find two numbers such that their sum multiplied by the sum of their squares is 65, and the difference multiplied by the difference of their square is 5.
Answer by mananth(16946)

(x+y)(x^2+y^2)=65
x(x^2+y^2)+y(x^2+y^2)=65
x^3+xy^2+x^2y+y^3=65..................(1)
(x-y)(x^2-y^2)=5
x(x^2-y^2)-y(x^2-y^2)=5
x^3-xy^2-x^2y+y^3=5.....................(2)
add (1) & (2)
2x^3+2y^3=70
x^3+y^3=35.........................(3)
subtract (2) from (1)
2x^2y+2xy^2=60
2xy(x+y)=60
xy(x+y)=30
from (3)
x^3+y^3=35.........................(3)
(x+y)^3-3xy(x+y)=35
(x+y)^3+3(30)=35
(x+y)^3=90+35
(x+y)^3=125
(x+y)=5...........................(4)
Now we know
(x-y)(x^2-y^2)=5
(x-y)(x+y)(x-y)=5
(x+y)(x-y)^2=5
5(x-y)^2=5
(x-y)^2=5/5=1
x-y=+/-1...................................(5)
solve (4)&(5)

we have (2,3) OR (3,2) as the solutions